### v1

# Paying to Do Better: Games with Payments between Learning Agents

Abstract
In repeated games, such as auctions, players typically use learning algorithms to choose their actions. The use of such autonomous learning agents has become widespread on online platforms. In this paper, we explore the impact of players incorporating monetary transfers into their agents’ algorithms, aiming to incentivize behavior in their favor. Our focus is on understanding when players have incentives to make use of monetary transfers, how these payments affect learning dynamics, and what the implications are for welfare and its distribution among the players. We propose a simple game-theoretic model to capture such scenarios. Our results on general games show that in a broad class of games, players benefit from letting their learning agents make payments to other learners during the game dynamics, and that in many cases, this kind of behavior improves welfare for all players. Our results on first- and second-price auctions show that in equilibria of the “payment policy game,” the agents’ dynamics can reach strong collusive outcomes with low revenue for the auctioneer. These results highlight a challenge for mechanism design in systems where automated learning agents can benefit from interacting with their peers outside the boundaries of the mechanism.
1 Introduction
Autonomous learning agents have become widespread on online platforms, playing an ever-larger role in many markets and economic ecosystems. One prominent example is the billion-dollar market of online ad auctions run by internet giants like Google, Meta, and Microsoft, in which, due to the speed of these auctions, bidding traffic is controlled by various automated bidding agents, provided either by the platforms themselves, or by a third-party. Typically, these agents get some high-level instructions from their users about their objectives and allowed action space, and then they interact with other agents in long sequences of repeated games (which could include thousands or millions of auctions per hour), using some learning algorithm to optimize the long-term payoffs for their users.
We are interested in repeated games of this sort where players use learning agents to play on their behalf. We take as a point of departure the observation that, in this setting, users may allow their agents to make payments to other agents in order to affect the game dynamic. In this paper, we study the effect of considering such payments between the agents on the long-term outcomes of the interaction.
The idea that payments and other types of financial benefits outside of the mechanism can affect behavior in the mechanism has been observed before, in areas ranging from blockchain fee markets, where it has become one of the guiding principles in the evolving design of these markets [^1] While research in this area has focused on manipulations involving the auctioneer (the “miner” who allocates space for transaction to be included in the next “block”), we study the scenario where payments take place among the bidding agents themselves. (see [24, 79] and [8, 9, 25]), to fair division [64, 65]. We defer a more extended discussion of related work to Appendix A.
The possibility that learning agents could transfer payments among themselves during their game dynamics raises several basic questions. First, when do players have incentives to let their agents use payments? Second, how do payments between the agents affect learning dynamics? And third, what are the long-term implications of payments among the agents on the players’ utilities, the social welfare, and in auctions, the revenue for the auctioneer?
As in much of the prior literature (e.g., [16, 28, 32, 38, 46, 53, 90]), we assume agents are no-regret learners; that is, they use learning strategies that result in outcomes satisfying the no-regret condition in the long term. Unfortunately, as is well known, in many games no-regret agents can end up playing strategies that yield low utility compared to cooperative outcomes that could potentially have been obtained in the game. E.g., in the prisoner’s dilemma, they end up defecting, since that is the dominant strategy. This has led to a great deal of interest in designing mechanisms where we can get, in some sense, better outcomes. Notably, in contrast to works in mechanism design in which an external entity provides incentives or recommendations aiming to improve welfare [7, 10, 62, 88], in our scenario there is no external entity, and players are only interested in their own individual payoffs.
Before giving an example, let us sketch the general structure of the model we have in mind. As mentioned, there is an underlying game and the model of what we call a payment policy game (or just a payment game) proceeds in two phases. In the first phase, each player $i$ chooses a payment policy, that determines, for each other agent $j$, how much agent $i$ will pay agent $j$ at each round in the agents’ dynamics as a function of the outcome (i.e., the joint action profile performed by the agents) in the underlying game in that round. In the second phase, the agents play the underlying game repeatedly, choosing their actions according to their learning algorithm (taking the payments into account). See Section 2 for further details.
Warm-up example: With this background, consider the variant of prisoner’s dilemma [^2] Note that this variant is somewhat nonstandard, in that the payoffs are not symmetric. shown in Figure 1. As mentioned above, standard analysis shows that when both players use no-regret agents, each agent will quickly learn to play its dominant strategy, which is to defect (i.e., play strategy $D$); the dynamics will converge to the Nash equilibrium of the stage game where both agents always defect and each player gets an expected payoff of $\nicefrac{{1}}{{3}}$.
{NiceTabular}
[…]
$C$ $D$
[…]
$D$ $2,0$ $\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}}$
Figure 1: A prisoner’s dilemma game.
Now we ask what can player $1$ achieve if he allows his agent to make payments to the other agent during the game dynamics? Consider the payment policy where player $1$’s agent pays player $2$’s agent $\nicefrac{{1}}{{3}}+\epsilon$ if player $2$’s agent cooperates, for some $\epsilon>0$, and pays nothing if player 2’s agent defects. It is easy to see that the game with these payments is dominance solvable: by repeatedly eliminating strictly dominated strategies, we are left only with the strategy $(D,C)$ as the unique equilibrium, and so the ensuing dynamics will converge to that outcome (see Appendix B), with payoffs $(2-\nicefrac{{1}}{{3}}-\epsilon,\nicefrac{{1}}{{3}}+\epsilon)$ for the two players.
It turns out that this outcome can be obtained as an ($\epsilon$-)equilibrium [^3] In an $\epsilon$-equilibrium, no player can increase their payoff by more than $\epsilon$ by altering their strategy. of the payment game. There is also a second $\epsilon$-equilibrium of the payment game, obtained by switching the roles of players $1$ and $2$, but due to the asymmetry in the payoffs of the underlying game, the social welfare in this second equilibrium is $1$, not $2$. Importantly, however, both equilibria are Pareto improvements over the Nash equilibrium of the underlying game, and the players’ welfare significantly increases compared to the game without payments.
We show, using a similar analysis, that we get these equilibria and improvement in social welfare in general prisoner’s dilemma games whenever the maximum welfare gap in the game is large (above a factor of $2$). Moreover, while in the game without payments the price of anarchy ($PoA$) and price of stability ($PoS$) —the ratio between the social welfare obtained from the worst-case (resp., best-case) Nash equilibrium and the optimal social welfare—are unbounded, with payments, both $PoS$ and $PoA$ are bounded by a factor of $2$ in the symmetric case, and in the asymmetric case $PoS$ is bounded, but $PoA$ is not, due to the welfare gap between the two equilibria of the payment game. See Appendix B.
Although prisoner’s dilemma is very simple (both players have strictly dominant strategies), the observations we made regarding this example apply more broadly. Specifically, in many games, players have incentives to use payments between their agents, and the use of payments can often enable players to reach more cooperative outcomes, with higher social welfare.
Our results: In this paper, we initiate the study of the incentives of players to use payments with their learning agents and the potential impact of such payments. We analyze first- and second-price auctions as our main case study in Section 3. In the following sections, we study general properties of payment games; Section 4 focuses on unilateral payments and Section 5 focuses on two-player games. Our results in all these settings provide comparative statics between the outcomes of learning dynamics without payments and those obtained in equilibria of the payment game.
More specifically, in second-price auctions, Theorem 1 shows that players have incentives to provide payment policies to their agents, and that in equilibrium of the payment game, the players may reach, in the long-term limit, full collusion where they capture almost the entire welfare, leaving the auctioneer with vanishing revenue. In first-price auctions, Theorem 2 shows that also here players have incentives to use payments, and equilibria of the payment game lead to low (but still positive) revenue. [^4] En route, we also provide the first simple dynamic approaching the minimum-revenue coarse correlated equilibrium of the first-price auction shown in [35]; see Observation 1. Then, in Section 4 (Theorems 3, and 4), we show that in a very broad class of finite games, [^5] Finite games are games with discrete action sets and bounded payoffs. Note that in the general model of auctions, the bid space is continuous, and so our analysis in this setting is different. there is at least some player who has the incentive to use payments, and we characterize cases where a payment policy by a single agent can lead the dynamics to the optimal welfare outcome while improving the payoff for the associated player as well. In Section 5, we show that in two-player games, payments always lead to Pareto improvements compared to the game without payments (Theorem 5), and that in a broad class of games, even including games with strictly dominant strategies (as our prisoner’s dilemma example shown above), at least one of the players would strictly benefit from using payments, and so not using payments is not stable (Theorem 6).
From a higher-level perspective, our results highlight a challenge for auction design and market design more broadly: we show that users of learning agents quite generally have incentives to allow their automated agents to make payments during their interactions. When these incentives are coupled with the right technology, particularly, sophisticated AI agents and flexible transaction media such as those on blockchain platforms, one can easily imagine an outlook where agents are trading among themselves “under the hood,” and the associated markets change their behavior, equilibria, and outcomes. Our results demonstrate that these changes can be very significant, underlining, on the one hand, the potential improvement in efficiency, but on the other hand, the need to better understand these changes for concrete markets, to be able to design them accordingly.
2 Model
In our model, we consider scenarios where we have players using learning agents in some repeated game, and study the setting where players can augment their agents by allowing them to make payments to other agents that depend on the actions of these agents during the game dynamics. The learning algorithms used by the agents are assumed to satisfy the regret-minimization property, [^6] We provide the definition of regret minimization and other standard definitions in Appendix C. but agents are not restricted to using any particular algorithm, and the players themselves are interested in their long-term payoffs. As our analysis relies on the regret-minimization property itself and considers the outcomes in the limit $T\rightarrow\infty$, where $T$ is the number of game rounds, our results apply regardless of the specific regret bounds of the learning algorithms used, as long as the agents achieve regret sublinear in $T$. Our model has the following components.
  • The game: There is an underlying game $\Gamma=\{[n],S,\{u_{i}\}_{i=1}^{n}\}$, where $[n]$ is the set of players, $S$ is the space of joint actions, and $u_{i}:S\rightarrow[0,1]$, $i=1,\dots,n$ are the utility functions in that game.
  • Agents: Every player uses a no-regret agent to play the game on their behalf for $T$ rounds. Denote the set of these agents by $A$, where $A_{i}\in A$ is the agent of player $i\in[n]$. Agents choose actions using their no-regret algorithms. The action of agent $A_{i}$ at time $t\in[T]$ is denoted $s_{i}^{t}$ and the action profile of all the agents is denoted $s^{t}$.
  • Payments: In addition to choosing actions, the agents make payments to other agents that depend on the actions $s^{t}$ chosen by the agents, according to payment policies defined by the associated players.
  • Payment policies: Each player $i$ chooses a policy that determines, for each action profile $s\in S$ and player $j\in[n]\setminus\{i\}$, the payment $p_{ij}(s)\in\mathbb{R}_{+}$ from agent $i$ to agent $j$ when the outcome in the most recent step is $s$. We assume that payments are bounded: $p_{ij}(s)\in[0,M]$, where $M\in\mathbb{R}_{+}$ is some large constant.
  • Agent utilities: In every step $t$, the agents play an action profile $s^{t}\in S$ and observe the realized outcome and payments. The utility for agent $i$ at time $t$ when the action profile that was played was $s^{t}$ is $v_{i}^{t}=u_{i}(s^{t})+\sum_{j\neq i}\big{(}p_{ji}(s^{t})-p_{ij}(s^{t})\big{)}$.
  • Player utilities: The players’ utilities are their long-term average payoffs: $U_{i}^{T}=\operatorname{\mathbb{E}}[\frac{1}{T}\sum_{t=1}^{T}v_{i}^{t}]$; we are particularly interested in the limit $U_{i}=\lim_{T\rightarrow\infty}U_{i}^{T}$.
  • Since the analysis of long-term payoffs ignores finite-time costs, we assume w.l.o.g. that if a player is asymptotically (in the limit $T\rightarrow\infty$) indifferent between making a payment and not making the payment, the player prefers not making the payment.
The model above defines a payment-policy game (in short, a payment game) between users of learning agents that is specified by $\Gamma$, $A$, and $T$. In order to understand the basic properties of payment-policy games and how the incentives of players change compared to games without payments, our analysis of equilibria and potential deviations focuses on one-shot, full-information payment-policy games.
A fundamental question that arises in our context is: “When would players have incentives to use payment policies?” Conversely, what are the conditions for a game to be “stable” in the sense that players do not have incentives to use such manipulations?
Definition 1.
A game $\Gamma$ is stable for a set of learning agents if zero payments are an equilibrium of the payment-policy game associated with $\Gamma$ with these agents. A game is stable if it is stable for any set of regret-minimizing agents for the players.
3 Auctions with Payments between Learning Agents
To get a deeper understanding of the incentives of users in online platforms to use payments with their learning agents and the potential impact of such payments, we focus on one class of games that have already been widely analyzed: auctions with automated bidders. We consider the simplest model of first- and second-price auctions, where at each time $t\in[T]$ a single identical item is sold. Each player $i$ has a value $v_{i}$ for the item, and we index the players in decreasing order of their values, $v_{1}\geq v_{2}\geq\dots\geq v_{n}$. At round $t$, each agent submits a bid $b_{i}^{t}$ and the auction mechanism determines the identity of the winner and a price $p^{t}$. The payoff for the winning agent from the auction is $v_{i}-p^{t}$; the other agents get payoff zero. In addition to payoffs from the auction, agents get a payoff according to the payments between the agents (as defined in Section 2). Utilities are additive over game rounds. In the first-price auction, the payment is equal to the highest bid; in the second-price auction, the payment is the second-highest bid. In both cases the highest bidder wins, and we assume tie-breaking according to the index of the agent (as is done, e.g., in [35]), so that if $i<j$ and $i$ and $j$ both have the same top bid, then $i$ is taken to be the winner. All our analyses are similar for arbitrary tie-breaking rules, but breaking ties in favor of the highest-value agent simplifies the presentation. Formal proofs for this section are in Appendix D.
Second-price auctions:
The second-price auction has been widely studied [86] (see Appendix A). While the auction is known to be incentive compatible, it is also known that truthful bidding forms only one out of many Nash equilibria of the stage game. In the context of learning dynamics, it has been shown that regret-minimizing agents in this auction do not converge to the truthful equilibrium; instead, they reach some degree of tacit collusion [53] with lower revenue to the auctioneer than that obtained under truthful bids.
Specifically, simplifying slightly, second-price auctions have two types of non-truthful equilibria. The first type is “overbidding equilibria,” where the high-value player bids anything above his value and the low-value player bids anything below the high value. The second type is “low-revenue equilibria,” where the high-value player bids anything above the low value and the low-value player bids anything below the low value. [^7] Both types of equilibria arise for any $n$; we describe the two-player case for ease of exposition. The latter type is especially interesting for our setting since any mixture of such equilibria is consistent with regret minimization, but yields high welfare to the players. This suggests that when learning agents exchange payments, the high-valued agent may be able to use payments to influence the dynamics with the other agents to induce a better equilibrium for himself.
We show that, indeed, when agents “trade outside of the mechanism” in this way during their dynamics, this kind of collusion emerges from the agents’ interaction in a strong way, potentially leading to zero revenue for the auctioneer.
Theorem 1.
The single-item second-price auction where the players’ top two values are $v_{1}>v_{2}$ is not stable for any regret-minimizing agents for the players, and the payment-policy game has an $\epsilon$-equilibrium where the players capture the full welfare and the auctioneer gets revenue zero.
Proof sketch: The idea of the proof is, first, to construct a unilateral payment policy for player $1$, parameterized by a single parameter $\epsilon$ that determines the amount that agent $1$ pays to the other agents, such that this policy induces a dynamic where player $1$’s (i.e., the high player’s) agent bids his value and all the other learning agents bid zero. We then show that for every other player, using the zero-payment policy is in fact a best response, and that the (net) payoff for the high-value player can be arbitrarily close to the full welfare, so player $1$ cannot improve his payoff by more than $\epsilon$ for any $\epsilon>0$. (Note, that with the second-price payment rule, if the agents were not restricted to using finite payments, the players would be locked into a game without an equilibrium, where each player tries to commit to making the highest bid and incentivize the others to bid zero.)
First-price auctions:
In contrast to the second-price auction, the first-price auction is known not to be incentive compatible. Intuitively, the direct dependence of the utility on the bid requires players to constantly optimize their responses to the bids of the other players. There has thus been significant interest in learning strategies, outcomes, and dynamics in this setting (see Appendix A for references). Interestingly, despite the significant difference from the truthful second-price auction, it is well known (and not hard to see) that the Nash equilibrium of the first-price auction yields the truthful outcome of the second-price auction. It has been shown that mean-based [19] (see Appendix C) regret-minimization dynamics [53] and best-response dynamics [71] can converge only to this outcome.
In [35], it was shown that the first-price auction also has “collusive” coarse correlated equilibria [^8] An equilibrium notion equivalent to all players having no regret. See [78] and Appendix C. (CCE) with lower revenue for the auctioneer than the second-price outcome, and higher utility for the players. The existence of such equilibria poses an opportunity for players to try and reach cooperative outcomes and increase their payoffs (at the expense of the auctioneer). However, for almost a decade, no natural process has been described, to the best of our knowledge, that attains any of these equilibria. Moreover, in [53], it is shown that even a scenario where players can manipulate the objective functions of their learning algorithms is not sufficient to reach any form of cooperation. In fact, convergence of standard no-regret agents to the second-price outcome induces a dominant strategy for users to provide the true objectives for their agents. To understand how cooperative outcomes could perhaps still be reached, we observe that in the collusive equilibria described in [35], the low-value player wins in some fraction of the auctions and has positive utility, whereas under simple no-regret dynamics (as those studied in [53]), only a single player wins in the limit. This suggests that payments between the agents during their dynamics could enable the agents to share the welfare and reach cooperative outcomes.
The following theorem shows that players have an incentive to use payments in a first-price auction, and in equilibria of the payment game, the players reach cooperative outcomes, with a significant reduction in revenue for the auctioneer.
Theorem 2.
The single-item first-price auction two players with values $v_{1}\geq v_{2}$ and any mean-based bidding agents is not stable, and every equilibrium of the payment game is a strong Pareto improvement (i.e., both players are better off) and has lower revenue for the auctioneer compared to the game without payments.
We provide here an outline of the proof, breaking it into a sequence of lemmas. The proofs of the lemmas are deferred to Appendix D. The first part of the proof is the construction of a unilateral payment policy for the high-value player and the analysis of the resulting dynamics. In this policy, the agent with the higher value pays $\eta>0$ to the low-value agent whenever the latter bids zero. We derive the (unique) mixed Nash equilibrium of the game with this payment policy, with $\eta$ as a parameter (Lemma 1). We then evaluate the utilities of the players in that equilibrium and show that the optimal value of $\eta$ for the high-value player is half the value of the second player (Lemma 2), and that in this case, the players have higher utilities than in the game without payments. We next prove that in the game between the agents, while the agents do not play a mixed equilibrium, and their strategies are correlated, nevertheless, the marginal distributions and payoffs are the same as in the mixed Nash equilibrium we derived (Lemma 3). Finally, using the fact that these utilities for the players are obtained with a unilateral payment by the high-value player, we show that these utilities bound from below the utilities in any equilibrium of the payment-policy game. We now state the lemmas carefully.
Lemma 1.
Consider the single-item first-price auction with two players and player values $v_{1}\geq v_{2}$, where player $1$’s payment policy is such that agent $1$ pays $\eta$ to agent $2$, where $0<\eta<v_{2}$, whenever agent $2$ bids zero. The resulting game between the agents has a unique Nash equilibrium where the cumulative density functions of the bids of agents $1$ and $2$ are $F_{1}(x)=\frac{\eta}{v_{2}-x}$ and $G_{2}(x)=\frac{v_{1}-v_{2}+\eta}{v_{1}-x},$ respectively (so the bid density functions are supported on $[0,v_{2}-\eta]$).
Lemma 2.
If the agents reach the Nash equilibrium utilities, the optimal value of $\eta$ for player $1$ is $v_{2}/2$. With this value of $\eta$, the utilities for the players are $v_{1}-v_{2}+\frac{v_{2}^{2}}{4v_{1}}$ for player $1$ and $\frac{v_{2}}{2}$ for player $2$. The winning frequency of player $2$ is in the interval $[0,3/8]$, with the lower bound being attained as $v_{2}/v_{1}\rightarrow 0$ and the upper bound being attained as $v_{2}/v_{1}\rightarrow 1$.
The result above is for Nash equilibrium. While in principle, no-regret agents may reach an arbitrary CCE and not necessarily this outcome, the next lemma shows that the family of Nash-equilibrium distributions parameterized by $\eta$ captures well the marginal distributions and utilities of the broad family of mean-based [19] no-regret agents in these games.
Lemma 3.
Consider the game where player $1$’s agent pays $\eta$ to player $2$’s agent whenever the latter agent bids zero. Fix a CCE with marginal distributions $F_{1}$ and $G_{2}$ to which the dynamics of mean-based agents converge with positive probability. Then, $F_{1}(b)=F_{1}^{NE}(b)=\frac{\eta}{v_{2}-b}$, $G_{2}(b)=G_{2}^{NE}(b)=\frac{v_{1}-v_{2}+\eta}{v_{1}-b}$, and the players’ utilities are the same as in the Nash equilibrium.
Remark.
Theorem 2 can be extended to some extent to $n$ players by a reduction to a game between the two top-valued players. In this reduction, all the lemmas extend with essentially no change in their proofs, so the result that first-price auctions are not stable extends as well. See Appendix D.2.1.
(a) Cumulative bid distributions
(b) Winning frequencies
Figure 2: Dynamics in first-price auctions.
Figure 2 shows simulation results with Hedge (a version of multiplicative weights [82]) agents in a sequence of $100$,$000$ auctions. The left panel compares the Nash equilibrium theoretical prediction from Lemma 3 with the empirical CDF from the agents’ dynamics with $v_{1}=1,v_{2}=0.5$. The right panel compares the theoretical winning frequencies of the players as specified in Lemma 2 with the empirical frequencies as a function of the ratio of values $v_{2}/v_{1}$ in multiple simulations. The long-term marginal bid distributions and winning frequencies are clearly consistent with the theoretical predictions. We observe similar results with other algorithm variants like follow the perturbed leader and linear multiplicative weights [3]. There is very little variation between simulation instances.
Interestingly, it turns out that these dynamics of the agents with payments recover, in a special case, a known collusive CCE distribution of the auction without payments.
Observation 1.
For symmetric bidders with $v_{1}=v_{2}=1$ and $\eta=1/e$, the Nash equilibrium bid distribution of the game with payment $\eta$ is the same as that of the minimum-revenue coarse correlated equilibrium of the standard first-price auction (without payments) given in [35].
Notice, however, that in our game, the players’ utilities are different from those of the CCE without payments. In particular, the payment $\eta$ that results in the this CCE distribution is not optimal for player $1$, so it is not an equilibrium of the payment-policy game.
We are now ready to conclude the proof of Theorem 2 by using the lemmas above to bound from below the utility for the high player in any Nash equilibrium of the payment game. The idea is to use the fact that the game analyzed in our proof is induced by a unilateral payment policy by the high-valued player. In any equilibrium, player $1$ can consider the deviation in which he rejects (pays back) all the payments he received and makes this unilateral payment, leading to a utility of $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$. The full proof is in Appendix D.2.
To evaluate the potential loss in welfare in the game considered above, we can analyze the winning frequencies from Lemma 2. The reduction in welfare is equal to the difference $v_{1}-v_{2}$ times the fraction of times in which the low-value player wins. This reduction in welfare is maximized, as can be verified by numerical calculation, when $v_{2}/v_{1}\approx 0.50959...$, at which point the reduction is approximately $8.96\%$. When considering the reduction in revenue, since the support of the bids is between zero and $v_{2}/2$, almost all the payments are strictly less than $v_{2}/2$, so the revenue is reduced to less than half the revenue of the outcome of the game without payments. Thus, the players manage to improve their utilities at the expense of the auctioneer, with a relatively small loss in efficiency.
4 Single-Player Manipulations
We now turn our attention to general games with $n$ players and focus on analyzing the effects of the simplest type of manipulations: payment policies where only a single agent pays. By making sufficiently high payments, clearly, an agent could induce any pure outcome in the game. The main questions are thus not about the power of payment policies, but about when a player has the incentive to use such policies, and what are the implications of such profitable payment policies for the welfare of other players. Proofs for this section are in Appendix E.
Notation: Denote the welfare of an outcome $s$ by $w(s)$ and the best-response utility of player $i$ to action profile $s$ by $u_{i}^{BR}(s)=\max_{s_{i}^{\prime}\in S_{i}}u_{i}(s_{i}^{\prime},s_{-i})$. For the following analysis, it is useful to define a notion of regret that measures the opportunity cost per step for player $i$ of playing the game where the strategy profile $x$ is played compared to the benchmark of his best action in hindsight in the (different) game where $y$ (rather than $x$) is played by the other players. To avoid confusion with standard regret (which considers the benchmark of the best response in hindsight to $x$), we call this “alternative regret.” Intuitively, it measures the extent to which a player regrets not incentivizing others to play $y$ rather than $x$.
Definition 2.
Let $x,y$ be two joint distributions on the players’ joint action space $S$. The alternative regret of player $i$ under the distribution of play $x$ compared to best responding to the alternative distribution of play $y$ is $R_{i}(x,y)=u_{i}^{BR}(y)-u_{i}(x)$.
At time $T$, we have an expected distribution of play $x(T)$ with welfare $w(x(T))$, where $x(T)$ approaches the set of CCEs of $\Gamma$ as $T\rightarrow\infty$. To simplify notation, we omit the dependence on $T$ and discuss it only when necessary. The optimal welfare is denoted $OPT=\max_{s\in S}w(s)$.
The following theorem characterizes the cases in which there is a player in the game who can gain from incentivizing the other agents to reach the optimal-welfare outcome.
Theorem 3.
Fix a finite game $\Gamma$ and a set $A$ of regret-minimizing agents. Let the optimal welfare outcome [^9] To simplify the analysis, we consider the generic case of a unique optimal outcome. In games with equalities in utility, this can be obtained by considering infinitesimal perturbations of the utilities. of $\Gamma$ be $y^{*}$. If $y^{*}$ is a Nash equilibrium, then there is a player $i$ such that $i$ increases his own payoff and the payoffs of all the other players by using a payment policy $p_{i}$ that induces $y^{*}$ as the unique long-term outcome of the agents’ dynamics. If $y^{*}$ is not a Nash equilibrium, the same holds if there exists a player $i$ for which the welfare gap $OPT-w(x)$ is greater than $\sum_{j\neq i}R_{j}(x,y^{*})$.
Analysis: To prove the theorem, we begin by establishing two lemmas that quantify the costs and gains for a player of pushing the agents’ dynamics to an arbitrary pure outcome.
To assure the convergence of arbitrary no-regret agents to $y^{*}$, the payments must induce $y^{*}$ is the unique CCE, [^10] With multiple CCEs, no-regret dynamics may fail to converge even in the time average [54]. which implies that it is also the unique Nash equilibrium. One way to do so is by making $y$ an equilibrium in dominant strategies. Although weaker requirements would be sufficient for our purpose (such as inducing a dominance-solvable game), using dominant strategies simplifies the analysis. Moreover, it turns out that it leads to the same payments and the same utilities for the players in the long term as inducing a unique CCE: [62] analyzed a mechanism-design scenario where a designer adds exogenous payments to the game, and they observed that the actual cost of inducing an outcome $s$ as an equilibrium in dominant strategies — which they called the optimal $k$-implementation of $s$, $k(s)$ — is equal to the cost of inducing that outcome as a Nash equilibrium. Combining this with the fact that an equilibrium in dominant strategies is also the unique CCE, we have that, indeed, the payments needed for inducing a unique CCE or an equilibrium in dominant strategies are the same. The intuition for this result is that further payments that are intended to assure dominant strategies are in fact made only off the equilibrium path, and are thus not actually made when the game is played. A similar effect was also used in [7] in the context of providing collateral contracts to mitigate strategic risk and incentivize investments.
This intuition also carries over, to some extent, to our setting, where one of the players uses a payment policy to manipulate the agents’ dynamics. But our case is different in several ways. First, importantly, in our case there are no external funds injected into the game. Instead, the payments that a player allows his agent to make to other agents are taken from his own payoffs. Second, as we show below, the cost of inducing an outcome $s$ as an equilibrium is lower than $k(s)$, since the player who wishes to induce $s$ as the outcome of the agents’ dynamics does not make payments to his own regret-minimizing agent. Third, since in our case the learning agents do not reason about payment policies but rather learn to respond over time, some agents will play dominated actions for some time, which will lead to additional short-term costs. This is formally expressed in the following lemma.
Lemma 4.
Fix a player $i$ and pure outcome $y$. Player $i$ can incentivize $y$ as the unique long-term outcome of the dynamic with an expected cost per time step of $\ cost_{i}(y)=k(y)-R_{i}(y,y)+o(1)$.
The next lemma specifies the condition under which, given a distribution $x$ to which the no-regret agents converge in expectation in $T$ rounds, there is a player who can use a payment policy to increase his payoff by pushing the dynamic to a different outcome.
Lemma 5.
Fix a finite game $\Gamma$ and a set of regret-minimizing agents for the players. If there exists a player $i$ and outcome $y$ such that $w(y)-w(x)>\sum_{j\neq i}R_{j}(x,y)$, then $i$ can increase his payoff by making payments that induce $y$ as the unique outcome of the agents’ dynamics.
Intuitively, the increase in welfare must be large enough to compensate the other players for their regret for playing $x$ relative to their best responses to the alternative outcome $y$.
Corollary 1.
If the dynamics without payments approach a welfare $w(x)$ lower than the welfare of some Nash equilibrium $y$, then, since in equilibrium $R_{i}(y,y)=0$, there is a player who can increase his payoff by using a payment policy that induces $y$ as the long-term outcome of the dynamics.
The lemmas above provide the basis for proving Theorem 3. See Appendix E for the proof.
From the analysis above, we observe that in many cases players can gain from using payment policies to divert the agents’ dynamics to more favorable outcomes for them. The following theorem shows that this is true for a broad class of games. For games with $PoS=1$, this is already implied by Theorem 3, but the following result is more general.
Theorem 4.
A finite game in which there exists a coarse correlated equilibrium with lower welfare than the best Nash equilibrium is not stable. This is true, in particular, if $PoA\neq PoS$.
Remark.
The converse is not true in general: there are games with $PoA=PoS$ that are not stable. One such example is the prisoner’s dilemma game in Figure 1.
5 Two-Player Games
In our prisoner’s dilemma example in Section 1, we saw that the players had incentives to use payments between their agents, and in equilibria of the payment game, both players had higher gains than in the game without payments. We now show that these results hold quite generally for two-player interactions. The proofs for this section are in Appendix F.
Theorem 5.
Every equilibrium of a two-player payment-policy game is a (possibly weak) Pareto improvement over the no-payments outcome of this game.
The intuition of the proof is that in order to have a positive payment in equilibrium, some player must benefit from making the payment while the other benefits from receiving it. If some player is worse off with the payments, then they always have a deviation that effectively cancels the payments. With more than two players, there may be players who cannot profitably counteract such payments made among others.
The next theorem addresses the question of incentives to use payments. The theorem considers games in which no-regret dynamics converge to a single CCE, and shows that despite the simplicity of learning in these games, in many cases players still have an incentive to use payment policies. The intuition for this result is that in two-player interactions, players can utilize payments to assume the role of a Stackelberg leader in the game.
Theorem 6.
Any finite two-player game with a unique coarse correlated equilibrium and two different pure-strategy Stackelberg outcomes (depending on which player is the Stackelberg leader) is not stable for any regret-minimizing agents for the players.
6 Conclusion
Autonomous learning agents are widely and increasingly used in many online economic interactions, such as auctions and other markets. Motivated by this transition to automated systems with sophisticated AI agents, we studied the scenario where automated agents also offer monetary transfers to other agents during their dynamics. Our results show that, generally, strategic users have incentives to use payments with their learning agents, and that this can have significant implications for the overall outcomes. Our findings highlight a challenge for mechanism design in the AI era, underlining the need to better understand both agent dynamics and user incentives in automated markets.
Appendix A Further Related Work
We follow a long research tradition on learning and dynamics in games, from early work of Brown, Robinson, Blackwell, and Hannan in the 1950s [13, 20, 43, 76], through seminal work in the following decades [39, 40, 44] and continuously since [1, 15, 26, 27, 37, 46, 51, 52, 53, 54, 55, 60, 83]. See [23, 45] for an overview of the foundations of the field. The notion of regret has been central throughout this work as a tool to define learning outcomes and objectives and design algorithms to achieve them.
No-regret learning has been broadly studied in auctions, with two prominent lines of work focusing on the price of anarchy in various auction formats with regret-minimizing bidders [14, 21, 28, 46, 77, 80, 84] and on the econometric problem of inferring bidder preferences [41, 67, 68, 69]. Other work has studied automated bidding algorithms for auctions with more complex features, such as budget constraints [11, 30, 37], or learning reserve prices [22, 61, 81]. In [2, 72], regret minimization is used as a prediction model for bidder behavior. Importantly, [67, 72] provide empirical evidence from large auction datasets showing that bidder behavior in practice is consistent with no-regret learning outcomes.
Learning dynamics in first- and second-price auctions and their convergence properties have been studied in a broad range of work [12, 18, 28, 29, 36, 53, 54, 71]. In [53, 54], the authors focus on analyzing the average outcomes of regret-minimization dynamics and their implications on user incentives. These papers consider a meta-game between the users of learning agents, which is induced by the long-term outcomes of the agents’ dynamics. This is close to our model of the payment game; however, payments have a very different impact on the agents’ dynamics and they lead to different analyses and results in auctions and other games.
The question of how distributed and self-interested players can reach cooperative or efficient outcomes is, of course, a very broad topic that has been studied in various (and largely separate) fields. The classic approach to this problem is from mechanism design [66, 70] and implementation theory [47, 59]. Somewhat closer to our work are [7, 62, 88], who study how an external mechanism designer can minimize the exogenous payments he needs to make to the players in order to implement particular outcomes in the game. By contrast, we study the impact of payments that automated agents could make among themselves; our analysis does not consider a mechanism-design problem [^11] Designing payment protocols to facilitate particular types of payments between the players to optimize social welfare is an interesting research avenue, but we do not pursue it here. and there are no external payments. An additional strand of work from the reinforcement learning (RL) and artificial intelligence literature studies how distributed RL algorithms can jointly reach cooperative outcomes in sequential social dilemmas [6, 33, 50, 58, 87]. A third line of work considers an approach for reaching cooperative outcomes called program equilibrium [57, 73, 74, 85], in which each player declares a program and each program can read the commitments made in the other programs and condition actions in the game on those commitments. These models, however, are very different from our model.
Finally, the idea of using payments to incentivize desirable actions is, of course, an old one, and represents one of the cornerstones of a rich literature on contract design and principal-agent problems. We refer the interested reader to [17, 31, 56] for an overview of this field. In most work in this area, there is a “principal” who provides incentives to the agents. As a result, he faces a contract-design or an optimization problem, rather than acting as a player in the game. Notable work that is closer to our scenario includes [48], who study multi-lateral payment contracts in one-shot interactions, [42, 89], who study learning in repeated contracts, and [75], who study scenarios where a player extracts fees from other players by using binding contracts. Our setting, however, differs significantly from the contracts setting, and our focus is on when players have incentives to augment learning agents with the ability to make payments, how payments among the agents affect their learning dynamics, and what the implications for the game outcomes are.
Appendix B Prisoner’s Dilemma Games
We begin by reviewing a known result (see, e.g., [49]) that is useful for our analysis.
Proposition 1.
In a dominance-solvable game, the empirical distribution of play of all types of regret-minimization dynamics converge with high probability to the unique Nash equilibrium.
Proof.
There are various ways to show this. Consider the following induction argument. The game has at least one order of iterated elimination of strictly dominated strategies. Index the players according to one such order, such that player $1$ has a dominant strategy, player $2$ has a dominant strategy in the sub-game where player $1$ plays his dominant strategy, and so on. Additionally, index the corresponding dominant strategies of the players using the same order, so that $a_{1}$ is the first action that is taken, $a_{2}$ is the second, and so on. Clearly, in any CCE, player $1$ plays his dominant strategy $a_{1}$; otherwise he has positive regret. Our induction step is simple: Fix any CCE. Given that players $11,...,k-$ 1 play actions $a_{1},...,a_{k-1}$, player $k$ has a strict best response to play $a_{k}$. By induction, this is true up to the last player $n$,so we have that our arbitrary CCE is the pure Nash equilibrium profile $(a_{1},...,a_{n})$. Therefore, regret-minimization dynamics converge to it with high probability. ∎
Equilibria of the game from Figure 1:
To approach the outcome described in the introduction as an equilibrium, we need to adjust the payment policy of player $1$ as follows. Player $1$’s agent pays player $2$’s agent the maximum payment whenever player $1$’s agent cooperates, and pays $\nicefrac{{1}}{{3}}+\epsilon$ when the outcome $(D,C)$ is obtained. Effectively, player $1$ blocks player $2$ from incentivizing player $1$’s agent to cooperate, so player $2$’s best response is not to pay anything and get $\nicefrac{{1}}{{3}}+\epsilon$. This is an $\epsilon$-equilibrium for all $\epsilon>0$, since player $1$ incentivizes outcome $(D,C)$ in the limit $T\rightarrow\infty$, and he cannot improve his long-term payoff by more than $\epsilon$. The other equilibrium, as mentioned, is the mirror image of the same equilibrium (replacing the payment policies of players $1$ and $2$).
{NiceTabular}
[…]
$C$ $D$
[…]
$D$ $\ x,0$ $1$,1\ 1 , 1
(a) The symmetric game: $x>y>1$.
{NiceTabular}
[…]
$C$ $D$
[…]
$D$ $\ x_{1},0$ $1$,1\ 1 , 1
(b) The asymmetric game: $x_{i}>y_{i}>1$.
PoS and PoA:
Consider the symmetric game in Figure 3(a), where in the prisoner’s dilemma we have $x>y>1$. The equilibria of the game depend of the welfare gap between the game outcomes. There are two cases. If $x\leq 2$, then the price of anarchy is trivially bounded by $2$, as this is the maximum utility gap in the game. If $x>2$, the payment game has two equilibria in which one agent pays the maximum amount when cooperating and pays $1+\epsilon$ when the other agent cooperates, and the other agent does not make any payments. Each of these equilibria yields a welfare of $x$. Since $x>y$, we have $\nicefrac{{2y}}{{x}}<2$, so the price of anarchy is at most $2$.
In the asymmetric game, we assume without loss of generality that $x_{1}\geq x_{2}$. We have three cases to analyze. If $x_{1},x_{2}\leq 2$, then the players do not have profitable payment policies, so, as before, $PoA=PoS$ and both are trivially bounded by $2$. If $x_{1}>2$ and $x_{2}\leq 2$, then only player $1$ has a profitable payment policy and the payment game has a single equilibrium with the outcome $(D,C)$, which yields a welfare of $x_{1}$ Since $x_{1}>y_{1},y_{2}$, we have $\frac{y_{1}+y_{2}}{x_{1}}<2$, $PoA=PoS$, and both bounded by $2$. Finally, if $x_{1},x_{2}>2$, as in the symmetric game, we have two equilibria for the payment game with outcomes $(D,C)$ or $(C,D)$. The $PoS$ in this case is at most $2$ by the same argument as in the previous case. However, the ratio $x_{1}/x_{2}$ between the two equilibria can be unbounded, so the $PoA$ can be unbounded.
Appendix C Additional Definitions
For completeness, we provide here several standard definitions.
Regret minimization:
For a given sequence of action profiles $s^{1},\dots,s^{T}$, the regret of agent $i$ is the difference in utility for $i$ between the actual utility in that sequence and the utility of the best fixed action in hindsight: $\text{Regret}_{i}(s^{1},\dots,s^{T})=\max_{s\in S_{i}}\big{[}\sum_{\tau=1}^{T}u_{i}(s,s_{-i}^{\tau})\big{]}-\sum_{\tau=1}^{T}u_{i}(s_{i}^{\tau},s_{-i}^{\tau}).$ A no-regret agent minimizes this quantity in the long term. There are several formulations for this property. We will use the following definition.
Definition 3.
An algorithm satisfies the (external) regret-minimization property if, for a time horizon parameter $T$ and any $T$-sequence of play of the other players $(s_{-i}^{1},\dots,s_{-i}^{T})$, where $s_{i}^{t}$ and $s_{-i}^{t}$ denote the actions taken at time $t$ by the algorithm and by the other players, respectively, we have that $\max_{s\in S_{i}}\big{[}\sum_{\tau=1}^{T}u_{i}(s,s_{-i}^{\tau})\big{]}-\sum_{\tau=1}^{T}u_{i}(s_{i}^{\tau},s_{-i}^{\tau})=o(T)$ with probability 1 in the limit as $T\rightarrow\infty$. An agent is regret-minimizing if it satisfies the regret-minimization property.
Coarse correlated equilibrium:
Coarse correlated equilibria are a weaker notion than correlated equilibria [5], also known as the Hannan set or Hannan consistent distributions [43]. The simplest definition for our purpose is the following.
Definition 4.
A joint distribution of play is a coarse correlated equilibrium if, under this distribution, all the players have in expectation regret at most zero.
Mean-based learning algorithms:
Mean-based learning algorithms [19] are a family of algorithms that play actions that are best responses to the history of play with high probability. This family was shown to include many standard no-regret algorithms, like multiplicative weights (see [3] and references therein), follow the perturbed leader [43, 51], and EXP3 [4].
Definition 5.
[19]. Let $\sigma_{i,t}=\sum_{t^{\prime}=1}^{t}u_{i,t^{\prime}}$, where $u_{i,t}$ is the utility of action $i$ at time $t$. An algorithm for the experts problem or the multi-armed bandits problem is $\gamma(T)$-mean-based if it is the case that whenever $\sigma_{i,t}<\sigma_{j,t}-\gamma(T)T$, then the probability that the algorithm pulls arm $i$ in round $t$ is at most $\gamma(T)$. An algorithm is mean-based if it is $\gamma(T)$-mean-based for some $\gamma(T)=o(1)$.
Appendix D Proofs for Section 3
D.1 Second-price auctions
Proof.
(Theorem 1). Consider a repeated second price auction with a single item sold in every step, and $n$ players with values $v_{1}>v_{2}\geq\dots\geq v_{n}$. Assume that the total payment that a player makes is bounded by some large constant $M>2nv_{1}$. Consider the following profile of payment policies. Every player $j>1$ uses the policy of always paying zero. Choose $\epsilon$ with $0<\epsilon<(v_{1}-v_{2})/n$. The payment policy for player $1$ is specified by the following conditions. (1) Whenever $b_{1}=v_{1}$, make a payment $p_{1j}$ to every player $j>1$ of $\epsilon/(n-1)$ if $b_{j}=0$ and zero otherwise. (2) If agent $1$ bids $b_{1}\neq v_{1}$, then he pays $p_{1j}=\frac{M}{n-1}$ to every other player $j>1$. Note that the maximum payment according to these conditions is $M$, as required.
Claim 1.
With these payments, bidding $v_{1}$ is a strictly dominant strategy for agent $1$.
Proof.
To see this, fix a bid profile for the other agents $2,\dots,n$, let $b_{0}$ denote the maximum bid of the other agents, and let $p$ denote the total payment to the other agents due to condition (1). We have the following cases.
Case 1 $b_{0}\leq v_{1}$: In this case, the bid $v_{1}$ gives agent $1$ a utility of $v_{1}-b_{0}-p$. Every other bid $b\neq v_{1}$ and $b\geq b_{0}$ gives a utility of $v_{1}-b_{0}-M$, which is strictly less. Every losing bid $b<b_{0}$ gives a utility of $-M$, which is also strictly less.
Case 2 $b_{0}>v_{1}$ (if agents overbid): Here every winning bid for agent $1$ gives a utility of $v_{1}-(b_{0}+M)$, every losing bid not equal to $v_{1}$ gives a utility of $-M$, and bidding exactly $v_{1}$ give a utility of $-p$, which is strictly higher. Therefore, bidding $v_{1}$ is a strictly dominant strategy, and agent $1$ who is minimizing regret will learn to only use this strategy. ∎
Next, we claim that since agent $1$ bids only $v_{1}$, every other agent has a strict best response, which is to bid zero. This is clear, since for every agent $j>1$, every winning bid gives negative utility, every losing bid $b\leq v_{1}$ and $b\neq 0$ gives zero utility, and bidding exactly zero give positive utility. Thus, with these payment policies, due to the regret-minimization property, agent $1$ always bids $v_{1}$ and wins the auction, while the other agents bid zero and get a payment of $\epsilon/(n-1)$ each. The utility for player $1$ in the long term is $v_{1}-\epsilon$, and the utility for every other player is $\epsilon/(n-1)$.
Claim 2.
The policy of always paying zero is a best response for every player $j>1$.
Proof.
We begin by looking at auctions with $n>2$. In every deviation by some agent $j>1$, in order to ensure that a bid $b<v_{1}$ is not dominated for agent $1$, agent $j$ must pay agent $1$ an amount of at least $M-v_{1}$. The utility for player $j$ in this case is bounded from above by $v_{j}-(M-v_{1})+\frac{M}{n-1}<2v_{1}-\frac{n-2}{n-1}M<2v_{1}\big{(}1-\frac{n(n-2)}{n-1}\big{)}<0$. That is, in every deviation by agent $j$ in which $j$ can win and get a positive utility from the auction (by incentivizing agent $1$ to bid below $v_{j}$ with some probability), the payment is too large, so the resulting utility is negative. So player $j$ prefers not to make payments and get a utility of $\epsilon/(n-1)$.
The case of $n=2$ requires a slightly different argument. In this case, agent $1$ pays agent $2$ an amount of $M$ whenever agent $1$ bids $b\neq v_{1}$. The fixed bid $v_{1}$ gives agent $1$ a utility of at least $v_{1}-\epsilon$. We claim that this implies that agent $1$ almost always wins the auction.
To see this, notice that a bid for which player $2$ pays agent $1$ an amount less than $M$ is still dominated and therefore not played by agent $1$. Suppose that there is a bid $b$ for which agent $2$ pays agent $1$ an amount of $M$. When agent $1$ loses with the bid $b$, he gets a utility of zero. When agent $1$ wins, he has an expected payoff of $v_{1}-\operatorname{\mathbb{E}}[b_{2}|b>b_{2}]$. Denote by $q$ the probability that $b$ is a winning bid for agent $1$. We have $v_{1}-\epsilon\leq q(v_{1}-\operatorname{\mathbb{E}}[b_{2}|b>b_{2}])\leq qv_{1},$ so we have $q\geq 1-\frac{\epsilon}{v_{1}}.$ Without loss of generality, we set $v_{1}=1$, and so agent $1$ wins at least $1-\epsilon$ fraction of the time.
Now, consider agent $2$. By the argument above, agent $2$’s payoff from using bids that are above zero is at most $v_{2}\cdot\epsilon<\epsilon$. Therefore, using such bids gives agent $2$ positive regret compared to bidding zero (when he bids zero, he gets a payment of $\epsilon$). Thus, agent $2$ almost always bids zero. It follows that player $2$ cannot make payments that increase his utility, so the policy of always paying zero is a best response. ∎
We are now almost done. With the policy profile that we have, player $1$ gets a utility of $v_{1}-\epsilon$. By taking a sufficiently small value of $\epsilon$, player $1$ gets a utility per time step that is arbitrarily close to $v_{1}$. Since $v_{1}$ is the optimal welfare in the game and player $1$ gets $\epsilon$ close to it for all $\epsilon>0$, this is an approximate best response for player $1$, and we have an $\epsilon$-Nash equilibrium. The auctioneer’s revenue in this equilibrium is zero, other than vanishing profits during the initial learning phase of the agents. ∎
D.2 First-price auctions
Proof.
(Lemma 1). We start by assuming that in the game between these agents, parameterized by $\eta$, there is a mixed Nash equilibrium with continuous CDFs, PDFs with full support on $[0,v_{2}-\eta]$, and player $2$ has a positive probability of bidding zero.
Consider player $1$’s utility. In equilibrium, he is indifferent between bidding zero and bidding $v_{2}-\eta$, giving us the size of the point mass for player $2$ at zero:
[…]
Player $1$’s utility when bidding $x$ is $u_{1}(x,y)=G_{2}(x)(v_{1}-x)-\chi(y)\eta$, where $\chi(y)$ is an indicator which equals $1$ if player $2$ bids zero (i.e., if $y=0$) and equals zero otherwise. Importantly, it is independent of $x$. Player $1$ is indifferent between all his bids, so $\frac{\partial u_{1}(x,y)}{\partial x}=0$, which gives
$G^{\prime}_{2}(x)(v_{1}-x)-G_{2}(x)=0.$
We have an ordinary differential equation with a unique solution of the form $G(x)=\frac{c}{v_{1}-x}.$ Given our initial assumption, we must have continuity at zero, so we get that
$c=v_{2}G_{2}(0)=v_{1}-v_{2}+\eta.$
Checking the consistency of the support (the CDF must be equal to $1$ at the top of the support):
$G_{2}(v_{2}-\eta)=\frac{v_{1}-v_{2}+\eta}{v_{1}-(v_{2}-\eta)}=1.$
Since player $2$ is indifferent between bidding $x>0$ and bidding zero and getting a utility of $\eta$, we have
$\eta=F_{1}(x)(v_{2}-x)\quad\Rightarrow\quad F_{1}(x)=\frac{\eta}{v_{2}-x}.$
Thus, we have a mixed Nash equilibrium with CDFs $F_{1}(x)=\frac{\eta}{v_{2}-x},\quad G_{2}(x)=\frac{v_{1}-v_{2}+\eta}{v_{1}-x}$, as stated in the lemma. ∎
Proof.
(Lemma 2). We start by computing the expected payoff of player $1$, which is given by
[…]
where the PDF is $f(x)=\frac{\eta}{(v_{2}-x)^{2}}$. From the distributions we calculated, we get that
$\operatorname{\mathbb{E}}[u_{1}]=v_{1}-v_{2}+\eta\Big{(}\frac{v_{2}-\eta}{v_{1}}\Big{)}.$
A straightforward calculation by differentiating this expression with respect to $\eta$ shows that the expectation is maximized when $\eta=v_{2}/2$, in which case its value is $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$, as claimed. With this choice of $\eta$, player 2’s utility is $v_{1}/2$.
The frequency at which the low agent (agent $2$) wins is given by:
[…]
If $v_{2}$ is small compared to $v_{1}$ (in the limit $v_{2}/v_{1}\rightarrow 0$), agent $2$ never wins. If $v_{2}$ is close to $v_{1}$ (in the limit $v_{2}/v_{1}\rightarrow 1$), agent $2$ will win $3/8$ of the time. ∎
Proof.
(Lemma 3) We begin by showing that the Nash equilibrium cumulative distributions bound the marginal distributions in our dynamics from below (i.e., the latter stochastically dominate the Nash equilibrium).
Claim 3.
The support of $F_{1}$ and $G_{2}$ is at most $v_{2}-\eta$, and for all $b$ in the support, we have that $F_{1}(b)\geq F_{1}^{NE}$, and $G_{2}(b)\geq G_{2}^{NE}(b)$.
Proof.
For agent $2$, all bids above $v_{2}-\eta$ are strictly dominated by bidding zero, and so are not in the support. For agent $1$, given the support of agent $2$, bids above $v_{2}-\eta$ give strictly less utility than bidding $v_{2}-\eta$, so a regret-minimizing agent will not play them.
For mean-based agents, we have that for all $b$ in the support, $u_{2}(b)=F_{1}(b)(v_{2}-b)\geq u_{2}(0)=\eta$ and $u_{1}(b)=G_{2}(b)(v_{1}-b)-\eta G_{2}(0)\geq v_{1}-v_{2}+\eta-\eta G_{2}(0)$. Hence we have $F_{1}(b)\geq\frac{\eta}{v_{2}-b}=F_{1}^{NE}$, and $G_{2}(b)\geq\frac{v_{1}-v_{2}+\eta}{v_{1}-b}=G_{2}^{NE}(b)$. ∎
Continuing the proof of Lemma 3, suppose by way of contradiction that $F_{1}(b^{\prime})>F_{1}^{NE}(b^{\prime})$ for some $b^{\prime}<v_{2}-\eta$. $F_{1}(b^{\prime})>F_{1}^{NE}(b^{\prime})$ implies $u_{2}(b^{\prime})=F_{1}(b^{\prime})(v_{2}-b^{\prime})>F_{1}^{NE}(b^{\prime})(v_{2}-b^{\prime})=\eta=u_{2}(0)$. Therefore, bidding zero is not in the support of player $2$.
However, if player $2$ never bids zero, then there are no payments in our CCE. Since the agents minimize regret, it must be a CCE of the standard first-price auction.
In the standard first-price auction, without payments, mean-based agents can converge only to CCEs that include bids above $v_{2}-\eta$ in their support (see [53]), a contradiction. Therefore, we have $F_{1}=F_{1}^{NE}$.
Next, suppose by way of contradiction that $G_{2}(b^{\prime})>G_{2}^{NE}(b^{\prime})$ for some $b^{\prime}<v_{2}-\eta$. This implies that $u_{1}(b^{\prime})=G_{2}(b^{\prime})(v_{1}-b^{\prime})-\eta G_{2}(0)>G_{2}^{NE}(b^{\prime})(v_{1}-b^{\prime})-\eta G_{2}(0)=v_{1}-v_{2}+\eta-\eta G_{2}(0)=u_{1}(v_{2}-\eta)$. That is, bidding $b^{\prime}$ gives player 1 a higher utility than bidding $v_{2}-\eta$, so the latter is not in the support of player 1, which is a contradiction, since the marginal distribution of player $1$ is $F_{1}$, which is supported on $[0,v_{2}-\eta]$. Thus, overall, we have $F_{1}=F_{1}^{NE}$ and $G_{2}=G_{2}^{NE}$. ∎
Proof.
(Theorem 2.) We will use the fact that the game analyzed in our proof is induced by a unilateral payment policy by the high-value player. In any equilibrium, player $1$ can consider the deviation in which he rejects (pays back) all the payments he received and makes this unilateral payment, leading to a utility of $v_{1}-v_{2}+v_{2}^{2}/4v_{1}$. Therefore, in any equilibrium, player $1$ gets at least this utility, which is higher than the utility of $v_{1}-v_{2}$ that he gets in the game without payments. As for the utility of the low-value player, since the high-value player’s utility is above the second price, player $1$ must use positive payments, otherwise the dynamics of mean-based agents could approach only the second-price outcome (as shown in [53]). Therefore, in the equilibrium of the payment-policy game, the low-value player has positive utility, an improvement over the game without payments.
The fact that the revenue decreases in any equilibrium is now straightforward. To see this explicitly, suppose by way of contradiction that the revenue $R$ is at least the revenue $v_{2}$ of the game without payments. The total welfare (of the players and the auctioneer) is $w=u_{1}+R+u_{2}\leq v_{1}$. We then get $v_{1}+\eta\Big{(}\frac{v_{1}-\eta}{v_{1}}\Big{)}+u_{2}\leq v_{1}$, which is a contradiction, since player $2$ cannot get negative utility in equilibrium.
By Lemma 2 and Lemma 3, player $1$ has a utility-improving deviation from the zero-payments profile for any mean-based regret-minimizing agents for the players, so the game is not stable for such agents; also, both players improve their utilities, and the revenue for the auctioneer decreases, as claimed. ∎
D.2.1 Extension of Lemma 3 to $n$ players.
To extend the lemmas in the proof of Theorem 2 to the case of $n$ players, consider the following analysis. Suppose we have $n$ players with values $v_{1}>v_{2}>v_{3}>\dots>v_{n}$, each using a regret-minimizing agent. If agent $2$ bids $v_{3}$, agent $1$ pays him $\eta>0$.
As in the proof of Lemma 1, we begin by assuming that in the game between these agents, parameterized by $\eta$, there is a mixed Nash equilibrium with continuous CDFs, PDFs with full support on $[v_{3},v_{2}-\eta]$, and player $2$ has a positive probability of bidding $v_{3}$.
We start by looking at player $1$’s utility. In equilibrium, he is indifferent between bidding $v_{3}$ and bidding $v_{2}-\eta$, giving us the (point-mass) probability that player $2$ bids $v_{3}$:
$v_{1}-(v_{2}-\eta)-G_{2}(v_{3})\eta=G_{2}(v_{3})(v_{1}-v_{3}-\eta)\quad\Rightarrow\quad G_{2}(v_{3})=\frac{v_{1}-v_{2}+\eta}{v_{1}-v_{3}}.$
Player $1$’s utility when bidding $x$ is $u_{1}(x,y)=G_{2}(x)(v_{1}-x)-\chi(y)\eta$, where $\chi(y)$ is an indicator which equals $1$ if player $2$ bids zero and equals zero otherwise. As before, note that it is independent of $x$. Player $1$ is indifferent between all his bids, so $\frac{\partial u_{1}(x,y)}{\partial x}=0$, which gives
$G^{\prime}_{2}(x)(v_{1}-x)-G_{2}(x)=0.$
We have an ordinary differential equation with a unique solution of the form
$G(x)=\frac{c}{v_{1}-x}.$
From continuity above $v_{3}$, we get that $c=v_{2}G_{2}(0)=v_{1}-v_{2}+\eta.$ Checking the consistency of the support (the CDF must be equal to $1$ at the top of the support):
$G_{2}(v_{2}-\eta)=\frac{v_{1}-v_{2}+\eta}{v_{1}-(v_{2}-\eta)}=1.$
Since player $2$ is indifferent between bidding $x>v_{3}$ and bidding $v_{3}$ and getting a utility of $\eta$, we have
$\eta=F_{1}(x)(v_{2}-x)\quad\Rightarrow\quad F_{1}(x)=\frac{\eta}{v_{2}-x}.$
So we have the CDFs
$F_{1}(x)=\frac{\eta}{v_{2}-x},\quad\quad G_{2}(x)=\frac{v_{1}-v_{2}+\eta}{v_{1}-x}.$
For players $i>2$, since there is a positive probability that both players $1$ and $2$ bid $v_{3}$ at the same time, any bid above $v_{3}$ wins the auction with positive probability and gives negative expected utility, whereas any bid $b\leq v_{3}$ gives a utility of zero. Therefore, any bid distribution for players $i>2$ that has zero weight on bids above $v_{3}$ forms a Nash equilibrium, together with the distributions $F_{1},G_{2}$ for the first two players.
The expected payoff of player $1$ is given by
[…]
where the PDF is $f(x)=\frac{\eta}{(v_{2}-x)^{2}}$. Thus, we have
[…]
So we see that in this equilibrium, player $1$ manages to increase his payoff with a unilateral payment. The optimal payment is $\eta^{*}=\min(v_{2}/2,v_{2}-v_{3})$. The payoff for player $2$ is $\eta$. Given the analysis above of Lemmas 1 and 2 with $n$ agents, the extension of Lemma 3 holds without any change to the proof, noticing the fact that since the minimum of the support is $v_{3}$, any player $j>2$ never wins the auction, so these agents can have arbitrary distributions with support at most $v_{3}$ with mean-based dynamics. Since this is a utility-improving deviation for player $1$ from the zero-utility payment profile, the auction is not stable with $n$ players.
Appendix E Proofs for Section 4
Proof.
(Lemma 4). We start constructing a payment policy for the agent of player $i$ by looking at the optimal $k$-implementation payments for inducing $y$ as a dominant strategy equilibrium; we will then adjust this payment profile to fit our setting in which there are no external payments and to minimize player $i$’s cost. A key point is that the payment amount $k(y)$ includes payments to agent $i$ that an external designer would have made for inducing $y_{i}$ as the dominant strategy for agent $i$. These payments are not feasible and not needed in our case. Instead, we change the payment policy of agent $i$ in the following way: for any action $s_{i}\neq y_{i}$, agent $i$ makes a payment of $\max_{s\in S}u_{i}(s)+1$, and distributes this payment equally to the other agents. These payments make all the other actions of agent $i$ strictly dominated by the action $y_{i}$.
Thus, to calculate the cost for player $i$, we need to subtract from the $k$-implementation payments the parts that are paid to player $i$, $R_{i}(y,y)$, and to add a learning-phase error term $\tilde{c}$ that results from the additional payments that agent $i$ makes when playing its dominated strategies, as well as from other times during the learning dynamics where the other agents play their dominated strategies.
Claim.
The expected learning-phase cost-per-time-step $\tilde{c}$ vanishes in the long term, that is, $\tilde{c}=o(1)$.
Proof.
$\tilde{c}$ is the expected total payment per time step that player $i$ makes when agents play strictly dominated strategies. Assume by way of contradiction that $\lim_{T\rightarrow\infty}\tilde{c}>0$, or that the limit does not exist. Both assumptions imply that the frequency at which dominated strategies are played in the agents’ dynamic is not vanishing, which contradicts the regret-minimization property of the agents. ∎
It follows that $cost_{i}(y)=k(y)-R_{i}(y,y)+o(1)$ as $T\rightarrow\infty$, as stated in the lemma. ∎
Proof.
(Lemma 5). We need to specify the requirement that player $i$ will increase his own utility after making the payments that induce $y$ as the unique unique outcome of the dynamics. That is, $cost_{i}(y)<u_{i}(y)-u_{i}(x)$. By Lemma 4 we have that for large enough $T$, player $i$ prefers making such payments if $u_{i}(y)-k(y)+R_{i}(y,y)>u_{i}(x)$. Rewriting the left terms we have
$u_{i}^{BR}(y)-k(y)>u_{i}(x)=w(x)-\sum_{j\neq i}u_{j}(x).$
Using the fact that, by definition [62], $k(y)=\big{(}\sum_{j}u_{j}^{BR}(y)\big{)}-w(y)$, we can rewrite the left-hand side as $w(y)-\sum_{j\neq i}u_{j}^{BR}(y).$ Rearranging the terms, we have
$w(y)-w(x)>\sum_{j\neq i}u_{j}^{BR}(y)-\sum_{j\neq i}u_{j}(x)=\sum_{j\neq i}R_{j}(x,y).$
Proof.
(Theorem 3). If the dynamics of the players’ regret-minimizing agents converge to the optimal-welfare outcome, $lim_{T\rightarrow\infty}x=y^{*}$, then there is no room for improvement and the theorem holds trivially (although weakly) with payments of zero. In the following, we assume for simplicity that $w(x)$ is bounded away [^12] There is also the scenario where $x=y^{*}$ for some values of $T$ but not in the limit. For such values of $T$ the theorem holds trivially as well, and so the assumption is without loss of generality. from $w(y^{*})$, that is, there exists $\epsilon>0$ and $T_{0}$ such that for all $T>T_{0}$, we have that $w(x)<w(y^{*})(1-\epsilon)$.
Suppose that $y^{*}$ is a Nash equilibrium. By the definition of Nash equilibrium, if some profile $s$ is a Nash equilibrium, then $R_{i}(s,s)=0$ for all $i$. Thus, by Lemma 4, any player can construct a payment policy that induces $y^{*}$ as the unique outcome of the agents’ dynamic with a cost-per-time-step of $o(1)$.
Therefore, to have a player $i$ who increases his payoff by inducing $y^{*}$ as the outcome of the dynamics, we only need to have that $u_{i}(y^{*})>u_{i}(x)$. Assume by way of contradiction that for all $i$, $u_{i}(x)\geq u_{i}(y^{*})$. Summing over the players, we have $\sum_{i}u_{i}(x)\geq\sum_{i}u_{i}(y^{*})$; that is, $w(x)\geq w(y^{*})$, a contradiction to the optimality of $y^{*}$. Thus, we have that for $T$ large enough that the learning-phase costs are small, there exists a player $i$ who can increase his payoff by inducing the optimal-welfare outcome as the outcome of the agents’ dynamics. If $y^{*}$ is a non-equilibrium outcome, the conditions and proof are given by Lemma 5 with $y=y^{*}$. Notice also that since under this payment policy $y^{*}$ dominates $x$ for all the agents, the utility for all the players is higher than in the game without payments. ∎
Proof.
(Theorem 4). The result follows from combining (the proof of) Lemma 5 with general properties of regret-minimization dynamics. Let $\Gamma$ be a game in which the Nash equilibrium with the highest welfare, denoted $\tilde{y}$, yields a welfare of $\tilde{w}$ (where $\tilde{w}$ is not necessarily equal to the optimal welfare $OPT$), and there exists a CCE $x$ with welfare $w(x)<\tilde{w}$.
It is known that for any CCE of a game, there are regret-minimizing agents that converge to that equilibrium [54, 63]. In particular, there are agents that converge arbitrarily close to $x$ [^13] Technically, the proofs of this convergence result use cycles of pure actions to approximate the time average of joint distributions, and the arbitrary approximation is due to the density of the rational numbers. We disregard this detail in our analysis. By Lemma 5 (in particular, Corollary 1), there exists a player who prefers using a payment policy that assures convergence to the highest welfare equilibrium $\tilde{y}$ compared to not making payments and reaching outcome $x$. Thus, we have a utility-improving deviation over the no-payments action profile and hence the game is not stable. ∎
Appendix F Proofs for Section 5
Proof.
(Theorem 5). The intuition for the proof in the two-player case is that to have a positive payment in equilibrium, some player must benefit from making the payment, while the other benefits from receiving the payment. With more than two players, there might be payments that are mutually beneficial to some subset of players but have negative externalities for others. In particular, with more than two players, there might be players who cannot profitably counteract payments made among others that do not benefit them.
Formally, consider a game $\Gamma$, a set $A$ of regret-minimizing learning agents for the players, and a time horizon $T$. Suppose, for the sake of contradiction, that there exists an equilibrium $p^{*}$ of the payment-policy game associated with ${\Gamma,A,T}$ where player $i$ has a lower utility than their expected utility over the $T$ rounds in the game without payments. Let $h$ denote the utility for player $i$ in the game without payments, and let $l$ denote the utility for player $i$ in the equilibrium $p^{*}$. By our assumption, we have $h>l$. Consider the following deviation from the payment policy $p^{*}$: player $i$ matches his payments to zero out all payments made by the other player and, additionally, reduces all of his other payments to zero. As a result, the two agents now observe in their dynamics the utilities of the original game $\Gamma$ without payments. Therefore, player $i$ has a utility of $h$. Since player $i$ has increased his utility from $l$ to $h$, we have a contradiction to $p^{*}$ being an equilibrium. Hence, the utilities of both players can only increase relative to the game without payments. ∎
Proof.
(Theorem 6). This result is related to a theorem in [54] that deals with games with dominant strategies in a meta-game in which users misreport their utility parameters to their learning agents. In our context of payments between the agents, the result is more general, applying to a larger family of games, including all dominance-solvable games, generic fully-mixed games, and socially-concave games [34] (a sub-class of concave games in which regret-minimization dynamics converge to Nash equilibria) with a unique equilibrium.
Consider as our starting point the action profile where the players use zero payments. In the game without payments, any regret-minimization dynamic will converge to the unique coarse correlated equilibrium (which is also the unique Nash equilibrium). To show that the game is not stable for any set of regret-minimizing agents for the players, we need to find a utility-improving deviation from the zero-payment profile for one of the two players.
We will use the fact that when the opponent does not make any payments, a player has the ability to alter his policy so as to commit to playing a pure action in the long term; this can be obtained by making large payments whenever the agent of this player plays different actions, inducing a dominant strategy for the agent. Since for at least one of the players there is a Stackelberg-equilibrium outcome that is different from the (unique) Nash-equilibrium outcome, there exists a player who obtains a higher payoff as a Stackelberg leader in the game, so this player prefers making such a commitment. Since the other agent is also minimizing regret, this agent will eventually learn to best reply to the fixed action of the first player’s agent, and the dynamics with these payments will converge to the Stackelberg outcome of the game. Thus, we have a utility-improving deviation from the zero-payment policy, so the game is not stable. Note that this is a different result from that of Theorem 4, since the Stackelberg outcome here is not a coarse correlated equilibrium of the game $\Gamma$. ∎
Acknowledgments
We thank Noam Nisan for his valuable feedback. Éva Tardos was supported in part by AFOSR grant FA9550-23-1-0410, AFOSR grant FA9550-231-0068. Joe Halpern was supported in part by AFOSR grant FA23862114029, MURI grant W911NF-19-1-0217, ARO grant W911NF-22-1-0061, NSF grant FMitF-2319186, and a grant from the Cooperative AI Foundation.